# Machine learning for option pricing:∑an empirical investigation of network architectures

### v2

@@ 6.2.3. Implied volatility
In Figure 6.33 the performance of the highway networks on the transformed implied volatility dataset is shown. On this dataset, we find similar relative MSEs and computational times as for the Black–Scholes and Heston pricing problems. In fact, the generalized highway network is again the superior network relative to the training time. However, just like in the Heston pricing problem scenario, the large MLP scores lower in absolute MSE. This indicates that the network performance on the implied volatility problem is also prone to larger network sizes. We expect to see performance improvements when training highway networks with more nodes and layers.
@@ 6.3.1. Black–Scholes model
The results of the DGM network performance on the Black–Scholes problem are visualized in Figure 6.35. We note that more operations must be computed inside a DGM network than in a highway network, resulting in slightly more parameters and hence in longer training time. This aligns with the numerical results.
Figure 6.35. The MLP, highway and DGM networks compared in terms of MSE on the test set (red) and training time (blue) for the Black–Scholes pricing problem. The small MLP contains $3$ layers with $50$ nodes and the large MLP contains $3$ layers with $500$ nodes.
The main difference between the highway layer and the DGM layer is that the original input $\mathbf{x}$ is used in each DGM layer, whereas the highway layer only depends on the output of the previous layer. The dependence on $\mathbf{x}$ introduces $4$ additional weight matrices per layer; see Figure 4.7. In Figure 6.35 no significant improvements due to this dependence compared to the generalized highway network can be observed. Indeed, the generalized highway network has a smaller MSE and requires less computational time for training. Possibly, the Black–Scholes problem is not complex enough for the additional non-linearity and recurrence in the DGM layers to cause improvements. When comparing the DGM network to the MLP and highway network in the training phase, cf. Figure 6.36, we notice that the DGM network consistently outperforms the highway and MLP networks on the training and validation sets. This observation may suggest that the DGM network is more flexible and can fit the training data better, possibly leading to overfitting, resulting in underperformance on the test set compared to the generalized highway network.
@@ 6.3.2. Heston model
In Figure 6.39 the performance of the DGM network alongside the smallest and largest $3$-layer MLPs and highway networks on the Heston pricing problem is shown. We notice that, unlike the Black–Scholes problem, the DGM network achieves a lower MSE than both the large MLP and the generalized highway network. The reason could be that the DGM has slightly more parameters which can lead to performance improvements; see also the discussion in Section 7.
Figure 6.39. The highway and DGM networks compared in terms of MSE on the test set (red) and training time (blue) for the Heston pricing problem.
Finally, in Figure 6.40 the results of the deep DGM and the no-recurrence DGM networks are shown. We notice that neither the no-recurrence network nor the deep DGM network significantly improve performance or training time. Interestingly, the no-recurrence DGM also requires slightly longer training time, despite a fewer amount of parameters in the network, compared to the DGM network. We can conclude that the DGM network outperforms both variations. In addition, we find that while the recurrence in the DGM network has decreased the MSE, the additional complexity it brings along does not make it preferable for option pricing problems in the tested setup.
@@ 6.3.3. Implied volatility
To complete the analysis, we evaluated the architectures on real implied volatility data obtained from the OptionMetrics database. Each daily implied volatility surface is given on a fixed grid of maturities
[…]
and put deltas
$\{-90,\,-80,\,-70,\,-60,\,-50,\,-40,\,-30,\,-20,\,-10\}.$
Each column of the dataset corresponds to one maturity–delta pair. We train the architectures with the same layers and nodes as in Table 6.42. For the MLP architecture, we consider the best one. Table 6.43 reports the results for all five architectures on the real implied volatility data. The performance on real implied volatility data is highly consistent with the synthetic experiments shown in Table 6.42. In both cases, the MSE falls in the range of $10^{-4}$ to $10^{-3}$. This indicates that the real dataset, although noisy, does not significantly increase the overall approximation difficulty compared with the synthetic implied-volatility surfaces. The ranking of architectures also follows similar patterns. The real-market experiment therefore provides an important robustness check: despite data imperfections, neural architectures achieve error levels comparable to the synthetic benchmark.
Model Layers Nodes Parameters Training Time (H) MSE
MLP 3 150 46,051 0.12 $3.44\cdot 10^{-4}$
Residual 3 50 5,501 0.12 $5.73\cdot 10^{-4}$
Highway 3 50 10,801 0.15 $3.80\cdot 10^{-4}$
Generalized Highway 3 50 16,101 0.13 $3.89\cdot 10^{-4}$
DGM 3 50 32,757 0.21 $4.64\cdot 10^{-4}$
It is also useful to compare neural-network approaches with classical non-neural methods for computing implied volatility. Numerical techniques such as Newton’s method, Jäckel’s algorithm, or Chebyshev interpolation can achieve extremely small absolute volatility errors (from about $10^{-12}$ to $10^{-6}$) when applied to well-behaved price inputs; see Glau et al. [15]. These methods are highly precise in controlled settings but do not directly approximate the entire volatility surface across strikes and maturities. Neural networks, by contrast, aim to learn the full surface in a single model, capturing complex patterns across both dimensions. As a result, the neural networks considered in this work reach mean absolute volatility errors around $10^{-2}$, reflecting the intrinsic difficulty of modelling a more irregular and high-dimensional function. This comparison highlights that while classical solvers provide a benchmark for precision, neural networks handle a broader approximation problem.
@@
Figure 6.44. The highway, MLP and DGM networks compared in terms of MSE on the test set (red) and training time (blue) for the transformed implied volatility problem.
@@
When we also include the results of the DGM variations, see Figure 6.45, we find that the no-recurrence DGM shows even better performance than the DGM network. This network architecture is similar to the generalized highway architecture, but contains one additional non-linear operation. In all the previous problems, the additional non-linear operation caused reductions in performance compared to the simpler generalized highway architecture. However, for the transformed implied volatility problem, we observe that the no-recurrence architecture shows a $51\%$ reduction in MSE compared to the DGM network, and a $66\%$ reduction compared to the generalized highway. Table 6.46 shows the MSEs, training times and parameters for all the networks of interest. Taking the training times into account, we can conclude that the no-recurrence DGM network shows the most favourable performance among the architectures considered for the implied volatility problem.
@@
Table 6.47. Comparison of DGM variants in terms of structural features, parameter count, training time, and MSE for B-S problem.
7. Capacity-normalized comparison of architectures
Although architectural comparisons across networks such as the Residual, Highway, Generalized Highway, and DGM networks are informative, differences in parameter counts can bias conclusions in favour of higher-capacity models. To isolate the effect of architectural design from raw model capacity, we conducted a capacity-normalized experiment in which all architectures were adjusted to contain approximately the same number of trainable parameters.
7.1. Experimental setup.
We designed new experiments where all architectures were resized to contain roughly 23,000 parameters. The following parameter-matched configurations were used:
@@ 7. Capacity-normalized comparison of architectures
All models were trained under identical conditions: Adam optimizer with learning rate $10^{-5}$, batch size of 64, and 200 epochs. The same training, validation, and test splits as in the main experiments were used (80% / 20% / 10%), and each experiment was repeated with three random seeds to report mean and standard deviation for both mean squared error (MSE) and training time.
7.2. Results.
Tables 7.48, 7.49, and 7.50 summarize the results for the Black–Scholes, Heston, and Implied Volatility problems, respectively.
@@
Table 7.48. Capacity-normalized comparison for the Black–Scholes problem (all networks $\sim$23k parameters).
@@
MLP 2 150 24,151 $0.98\pm 0.02$ $(7.30\pm 0.53)\cdot 10^{-6}$
Residual 3 104 23,713 $0.98\pm 0.02$ $(1.53\pm 0.14)\cdot 10^{-5}$
Highway 4 59 22,834 $0.18\pm 0.05$ $(2.63\pm 0.27)\cdot 10^{-5}$
Generic Highway 3 59 23,365 $0.34\pm 0.03$ $(2.44\pm 0.04)\cdot 10^{-5}$
DGM 2 50 24,467 $0.65\pm 0.47$ $(2.33\pm 0.27)\cdot 10^{-5}$
Table 7.49. Capacity-normalized comparison for the Heston pricing problem (all networks $\sim$23k parameters).
@@
Table 7.50. Capacity-normalized comparison for the implied volatility problem (all networks $\sim$23k parameters).
7.3. Discussion.
The capacity-normalized results show that when the model size is kept constant, architectural differences lead to substantial variation in performance across the three problems. In the Black–Scholes setting, Highway-type architectures achieve the best accuracy, indicating that gating mechanisms are well suited for smooth PDE solutions. For the more complex Heston model, however, the simple MLP yields the lowest error, suggesting that the inductive biases of gated or residual architectures may be overly restrictive under a fixed parameter budget. One should also notice though that the highway and generalized highway are significantly faster than the MLP and residual network solutions. In the implied-volatility problem, DGM and MLP architectures perform similarly and outperform residual-type models.
Our results show that there is no single best neural network architecture for all problems: each type of network works well for some tasks but not for others.
Importantly, even when all models have the same number of parameters, the differences in performance did not disappear. This means that the results are not caused by some models being bigger or having more capacity. Instead, the differences come from the way each architecture is built.
@@ 8. Conclusion
The empirical results obtained in this work allow us to draw several practical conclusions regarding the choice of neural network architectures for option pricing and implied volatility estimation:
  • Across the Black–Scholes, Heston, and transformed implied volatility problems, we observe that increasing the number of parameters generally improves accuracy when combined with an appropriate architecture. Larger models typically achieve lower MSEs, although the rate of improvement depends strongly on the underlying problem.
  • We did not find evidence that losses during training were a good indicator for the performance of the network during evaluation. Instead, we found that networks performing well on the training data set did in no case outperform the other networks on the testing data (see e.g. the DGM network on the Black–Scholes pricing problem; Figures 6.35 and 6.36). We would have liked to see such a relationship, as it allows us to terminate the training cycle early when no improvement is made on the training set compared to other networks, saving valuable computational time.
  • During training, the validation loss was often lower than the training loss. This indicates that we could possibly improve the accuracy of the networks by training them for more epochs. In Liu et al. [22], the results on the Heston pricing problem were more accurate, training for $3000$ cycles compared to $200$ in this work, while using an identical configuration for the MLPs.
  • For the transformed implied volatility problem, the generalized highway network also performed well, but was outperformed by the no-recurrence DGM architecture. Interestingly, this architecture did not perform competitively on the pricing problems, highlighting that architectural benefits are task-dependent.
  • In the case of the transformed implied volatility problem, we also found good performance results for the generalized highway network, which was however outperformed by the no-recurrence DGM architecture. This architecture is similar to the generalized highway network, but did not yield good results on the pricing problems.
  • When training networks on the implied volatility dataset, we found poor performance for all networks, and no improvements when using more complex architectures. When transforming the dataset slightly, as explained in Section 5.2, the prediction accuracy increased significantly, and we found similar results in network architecture performance compared to the Black–Scholes and Heston pricing problems. This observation suggests that the training of networks is highly dependent on the formulation and the dataset. When training neural networks, we should try to ensure that gradients are bounded. A possible way to do this is to transform the input variables so that the gradient with respect to the input variables does not approach zero or infinity.
@@
  • The evaluation on real implied volatility data shows that the architectures maintain error magnitudes and relative performance patterns comparable to the synthetic experiment, indicating that the task-dependent behaviours observed throughout the study persist under real-market conditions.
  • Finally, in Section 7 we supplemented our analysis with a capacity-normalized comparison in which all architectures were matched to approximately the same number of parameters ($\approx$23k). These results demonstrate that no architecture is universally superior. Instead, performance depends strongly on the underlying task: the generalized highway network performs best for the Black–Scholes problem, the MLP performs best for the Heston model, and the DGM or MLP architectures perform best for the implied volatility mapping. This comparison confirms that our earlier conclusions were not artifacts of differing model sizes and highlights that architectural advantages are inherently problem-specific.
@@ Appendix A Summary of network architectures
MLP 2, 3 50–500 ReLU Glorot 2,851-504,001 6.18, 6.20, 6.23, 6.25
Residual 3 50 $\tanh$ Glorot 7,951 6.29, 6.31, Fig. 6.32
Highway 3 50 $\tanh$ Glorot 15,601 6.29, 6.31, Fig. 6.32
Generalized Highway 3 50 $\tanh$ Glorot 23,251 6.29, 6.31, Fig. 6.32
Table A.51. Overview of feedforward neural network architectures and corresponding hyperparameter configurations.
Architecture Layers Nodes Activation Initializer Parameters Results tables
DGM 3 50 $\tanh$ Glorot Normal 33,459 6.38,6.42,6.46
Deep DGM 5 50 $\tanh$ Glorot Normal 49,959 6.38,6.42,6.46
No-Recurrence DGM 3 50 $\tanh$ Glorot Normal 31,059 6.38,6.42,6.46

### v3

@@ 6.2.3. Implied volatility
In Figure 6.19 the performance of the highway networks on the transformed implied volatility dataset is shown. On this dataset, we find similar relative MSEs and computational times as for the Black–Scholes and Heston pricing problems. In fact, the generalized highway network is again the superior network relative to the training time. However, just like in the Heston pricing problem scenario, the large MLP scores lower in absolute MSE. This indicates that the network performance on the implied volatility problem is sensitive to larger network sizes. We expect to see performance improvements when training highway networks with more nodes and layers.
@@ 6.3.1. Black–Scholes model
The results of the DGM network performance in the Black–Scholes problem are visualized in Figure 6.21. We note that more operations must be computed inside a DGM network than in a highway network, resulting in slightly more parameters and hence in longer training time. This aligns with the numerical results.
Figure 6.21. The MLP, highway and DGM networks compared in terms of MSE on the test set (red) and training time (blue) for the Black–Scholes model. The small MLP contains $3$ layers with $50$ nodes and the large MLP contains $3$ layers with $500$ nodes.
The main difference between the highway layer and the DGM layer is that the original input $\mathbf{x}$ is used in each DGM layer, whereas the highway layer only depends on the output of the previous layer. The dependence on $\mathbf{x}$ introduces $4$ additional weight matrices per layer; see Figure 4.5. In Figure 6.21 no significant improvements due to this dependence compared to the generalized highway network can be observed. Indeed, the generalized highway network has a smaller MSE and requires less computational time for training. Possibly, the Black–Scholes problem is not complex enough for the additional non-linearity and recurrence in the DGM layers to cause improvements.
Figure 6.22 reports the training and validation MSE for the DGM, generalised highway, and large MLP architectures applied to the Black–Scholes problem. The training and validation losses decrease concurrently for all models throughout the optimization and remain closely aligned, with no evidence of a systematic divergence between the two curves. This indicates that none of the architectures exhibit overfitting over the training horizon considered. Although the validation losses display higher variance, particularly at later epochs, their overall downward trend mirrors that of the training losses, which is consistent with stochastic simulation and Monte Carlo sampling effects rather than deteriorating generalization. Among the three architectures, DGM attains the lowest error levels, followed by the generalized highway network and the large MLP, reflecting differences in architectural inductive bias rather than generalization failure. Overall, the results demonstrate stable training dynamics and good generalization performance across all models.
@@ 6.3.2. Heston model
In Figure 6.25 the performance of the DGM network alongside the smallest and largest $3$-layer MLPs and highway networks in the Heston pricing problem is shown. We notice that, unlike the Black–Scholes problem, the DGM network achieves a lower MSE than both the large MLP and the generalized highway network. The reason could be that the DGM has slightly more parameters which can lead to performance improvements; see also the discussion in Section 6.4.
Moreover, in the same figure (Fig. 6.25) the results of the deep DGM and the no-recurrence DGM networks are included. We notice that neither the no-recurrence network nor the deep DGM network significantly improve performance or training time. Interestingly, the no-recurrence DGM also requires slightly longer training time, despite a fewer number of parameters in the network, compared to the DGM network. We can conclude that the DGM network outperforms both variations. In addition, we find that while the recurrence in the DGM network has decreased the MSE, the additional complexity it brings along does not make it preferable for option pricing problems in the tested setup.
@@ 6.3.3. Implied volatility
When we also include the results of the DGM variations, see again Figure 6.28, we find that the no-recurrence DGM shows even better performance than the DGM network. This network architecture is similar to the generalized highway architecture, but contains one additional non-linear operation. In all the previous problems, the additional non-linear operation caused reductions in performance compared to the simpler generalized highway architecture. However, for the transformed implied volatility problem, we observe that the no-recurrence architecture shows a $51\%$ reduction in MSE compared to the DGM network, and a $66\%$ reduction compared to the generalized highway. Table 6.29 shows the MSEs, training times and parameters for all the networks of interest. Taking the training times into account, we can conclude that the no-recurrence DGM network shows the most favourable performance among the architectures considered for the implied volatility problem.
@@
Table 6.30. Comparison of DGM variants in terms of structural features, parameter count, training time, and MSE for the Black-Scholes problem.
6.4. Capacity-normalized comparison of architectures
Although architectural comparisons across networks such as the residual, highway, generalized highway, and DGM networks are informative, the differences in parameter counts can bias conclusions in favour of higher-capacity models. Therefore, in order to isolate the effect of the architectural design from the raw model capacity, we conducted a capacity-normalized experiment in which all architectures were adjusted to contain approximately the same number of trainable parameters.
6.4.1. Experimental setup.
We designed new experiments in which all architectures were resized to contain roughly the same number of parameters. The following parameter-matched configurations were used:
@@ 6.4. Capacity-normalized comparison of architectures
All models were trained under identical conditions: Adam optimizer with learning rate $10^{-5}$, batch size of 64, and 200 epochs. The same training, validation, and test splits as in the main experiments were used (i.e. 80% / 20% / 10%), and each experiment was repeated with three random seeds to report mean and standard deviation for both mean squared error (MSE) and training time. Tables 6.31, 6.32, and 6.33 summarize the results for the Black–Scholes, Heston, and implied volatility problems, respectively.
@@
Table 6.31. Capacity-normalized comparison for the Black–Scholes problem (all networks have approximately 23k parameters).
@@
MLP 2 150 24,151 $0.92\pm 0.05$ $(6.2\pm 5.0)\cdot 10^{-6}$
Residual 3 104 23,713 $0.95\pm 0.02$ $(1.48\pm 0.05)\cdot 10^{-5}$
Highway 4 59 22,834 $1.04\pm 0.04$ $(2.81\pm 0.35)\cdot 10^{-5}$
Generic Highway 3 59 23,365 $1.06\pm 0.05$ $(2.31\pm 0.18)\cdot 10^{-5}$
DGM 2 50 24,467 $1.10\pm 0.02$ $(3.31\pm 0.16)\cdot 10^{-5}$
Table 6.32. Capacity-normalized comparison for the Heston problem (all networks have approximately 23k parameters).
@@
Table 6.33. Capacity-normalized comparison for the implied volatility problem (all networks have approximately 23k parameters).
6.4.2. Discussion.
The capacity-normalized results show, first of all, that the training time is determined by the number of parameters and not the architecture. Indeed, as expected, all models in the capacity-normalized experiment need approximately the same time to be trained. However, we still observe that architectural differences lead to substantial variation in performance across the three problems. In the Black–Scholes setting, highway-type architectures achieve the highest accuracy, indicating that gating mechanisms are well suited for this problem; this is largely consistent with the results of the previous subsections (see Fig. 6.13). On the other hand, for the more complex Heston model, the simple MLP yields the lowest error, suggesting that the inductive biases of gated or residual architectures may be overly restrictive under a fixed parameter budget. In addition, for the implied-volatility problem, DGM and MLP architectures perform similarly and outperform residual-type models.
Our results show that there is no single best neural network architecture for all problems: each type of network works well for some tasks but not for others. Importantly, even when all models have the same number of parameters, the differences in performance did not disappear. This means that the results are not caused by some models being bigger or having more capacity. Instead, the differences come from the way each architecture is built.
6.5. Application to real data and comparison with other methods
In order to complete the analysis, we evaluated the architectures on real implied volatility data obtained from the OptionMetrics database. Each daily implied volatility surface is given on a fixed grid of maturities
[…]
and put deltas
$\{-90,\,-80,\,-70,\,-60,\,-50,\,-40,\,-30,\,-20,\,-10\}.$
Each column of the dataset corresponds to one maturity–delta pair. We train the architectures with the same layers and nodes as in Table 6.27; for the MLP architecture, we consider the best one. Table 6.34 reports the results for all five architectures on the real implied volatility data. The performance on real implied volatility data is highly consistent with the synthetic experiments shown in Table 6.27. In both cases, the MSE falls in the range of $10^{-3}$ to $10^{-4}$. This indicates that the real dataset, although noisy, does not significantly increase the overall approximation difficulty compared to the synthetic implied-volatility surfaces. The ranking of architectures also follows similar patterns. The real-market experiment therefore provides an important robustness check: despite data imperfections, neural architectures achieve error levels comparable to the synthetic benchmark.
Model Layers Nodes Parameters Training Time (H) MSE
MLP 3 150 46,051 0.12 $3.44\cdot 10^{-4}$
Residual 3 50 5,501 0.12 $5.73\cdot 10^{-4}$
Highway 3 50 10,801 0.15 $3.80\cdot 10^{-4}$
Generalized Highway 3 50 16,101 0.13 $3.89\cdot 10^{-4}$
DGM 3 50 32,757 0.21 $4.64\cdot 10^{-4}$
Finally, it is useful to compare the neural-network approaches with classical non-learning methods for computing implied volatility. Numerical techniques such as Jäckel’s algorithm, or Chebyshev interpolation can achieve extremely small absolute volatility errors (from about $10^{-6}$ to $10^{-12}$) when applied to well-behaved price inputs; see Glau et al. [15]. These methods are highly precise in controlled settings but do not directly approximate the entire volatility surface across strikes and maturities; instead they apply different schemes depending on the regime of the input (e.g. deep in-the-money, etc). Neural networks, by contrast, aim to learn the full surface in a single model, capturing complex patterns across both dimensions. As a result, the neural networks considered in this work reach mean absolute volatility errors around $10^{-2}$, reflecting the intrinsic difficulty of modelling a more irregular and high-dimensional function. This comparison highlights that, while classical solvers provide a benchmark for precision, neural networks handle a broader approximation problem.
@@ 7. Conclusion
The empirical results obtained in this work allow us to draw several practical conclusions regarding the choice of neural network architectures for option pricing and implied volatility computation:
  • Across the Black–Scholes, Heston, and transformed implied volatility problems, we observe that increasing the number of parameters generally improves accuracy when combined with an appropriate architecture. Larger models typically achieve lower MSEs, although the rate of improvement depends strongly on the underlying problem. However, they also require more training time.
  • Across all architectures, training and validation losses evolve consistently, indicating stable optimization and good generalization without evidence of overfitting. Although validation losses exhibit higher variance, their overall behaviour closely follows the training performance.
  • The validation loss was often lower during training than the training loss. This indicates that we could improve the accuracy of the networks by training them for more epochs. In Liu et al. [22], the results on the Heston pricing problem were more accurate, training for $3000$ cycles compared to $200$ in this work, while using an identical configuration for the MLPs.
  • We observe that highway-type network architectures outperformed the other networks for the Black–Scholes problem, when we compared the MSE relative to the training time. This was consistent also in the capacity-normalized experiment. The same observation is true for the Heston model, but was not consistently observed in the capacity-normalized experiment.
  • In the transformed implied volatility problem, the generalized highway network also performed well, but was outperformed by the no-recurrence DGM architecture. Interestingly, this architecture did not perform competitively on the pricing problems, highlighting that architectural benefits are task-dependent.
  • When training networks on the implied volatility dataset, we found poor performance for all networks, and no improvements when using more complex architectures. When transforming the dataset as explained in Section 5.2, the prediction accuracy increased significantly, and we found similar results in network architecture performance compared to the Black–Scholes and Heston pricing problems. This observation suggests that the training of networks is highly dependent on the formulation and the dataset. When training neural networks, we should try to ensure that gradients are bounded. A possible way to do this is to transform the input variables so that the gradients with respect to them do not approach zero or infinity.
@@
  • The evaluation on real implied volatility data shows that the architectures maintain error magnitudes and relative performance patterns comparable to the synthetic data experiments, indicating that the task-dependent behaviours observed throughout the study persist under real-market conditions.
@@ Appendix A Summary of network architectures
MLP 2, 3 50–500 ReLU Glorot 2,851-504,001 6.4, 6.6, 6.9, 6.11
Residual 3 50 $\tanh$ Glorot 7,951 6.15, 6.17, 6.18
Highway 3 50 $\tanh$ Glorot 15,601 6.15, 6.17, 6.18
Generalized Highway 3 50 $\tanh$ Glorot 23,251 6.15, 6.17, 6.18
Table A.1. Overview of feedforward and highway neural network architectures and corresponding hyperparameter configurations.
DGM 3 50 $\tanh$ Glorot Normal 33,459 6.24,6.27,6.29
Deep DGM 5 50 $\tanh$ Glorot Normal 49,959 6.24,6.27,6.29
No-Recurrence DGM 3 50 $\tanh$ Glorot Normal 31,059 6.24,6.27,6.29